# Probing quantum anomaly corrections on black hole physics through chaos

### v1

@@ Abstract
The black hole horizon can induce chaotic motion of particles around the black hole. The original integrable motion of particles can transit to the chaotic motion when approaching black hole horizon. In this work, we consider the black hole background where quantum conformal anomaly correction is taken into account. We use the Poincare section and Lyapunov exponent as representative probes to illustrate chaos. We investigate the effect of anomaly coefficient on the chaos of particles and find that quantum anomaly enhances the chaos of particle motions. As the chaotic orbits of particles can leave an imprint on the gravitational wave signals of the extreme mass ratio inspirals, our results pave the way to detect conformal anomaly effect in actual observations.
@@ 1 Introduction:
Besides the dynamical properties of particles around black hole, black hole itself also has interesting dynamical properties. There are four law of black hole dynamics which is in close analogy with thermodynamical four laws [6]. The discovery of Hawking radiation [7] promotes the four laws of black hole dynamics to black hole thermodynamics and black hole behaves like an ordinary thermodynamical system with its entropy proportional to the horizon area [8]. Moreover, if the cosmological constant can be treated as the thermodynamical pressure [9], the equation of state of RN-AdS black hole is the same as the equation of state in van der-Waals liquid-gas system, thus by analogy, the black hole can exhibit rich phase transition structures [10, 11]. For RN-AdS black hole, there exists a critical point where the phase transition becomes second order and the critical exponents coincide with the mean field theory result [10].
@@
The chaotic motion of particles moving around black hole receives lots of attentions, there are many investigations from different point of views. Ref.[18, 19, 20, 21, 22, 23, 24, 25] investigate the chaos bound and its possible violations. Ref.[26, 27, 28] investigates the relation between chaos and black hole phase transition. Ref.[29, 30] investigate if the chaotic motions around black hole can be used to probe the dark matter in the universe. While Ref.[31, 14] considered the chaotic particle motions in modified gravity theories inspired by quantum effects. Moreover, recently Ref.[32] considered the imprint of particles’ chaotic motion on the gravitational waves for extreme mass ratio inspirals (EMRI) system.
@@ 3 Poincare section of conformal anomaly corrected black hole:
with the blackening factor corrected by the conformal anomaly given in Eq.(5). By redefining $E\to E/m$ and $p\to p/m$, the mass $m$ can be set to unity without loss of generality.
In order to numerically solve the above equations of motion, we choose the following parameter values
@@
The initial value of $r$, $\theta$ and $p_{r}$ is chosen randomly while the initial value of $p_{\theta}$ is solved by the energy constraint Eq.(13). In Fig.1, we show the Poincare section of the particle trajectories which is projected in the $(r,p_{r})$ plane. The section is defined by the constraint condition $\theta=0$ and $p_{\theta}>0$. We gradually increase the energy to see the change of Poincare sections. It can be found that under the influence of conformal anomaly $\alpha_{c}$, by increasing the energy which corresponds to approaching the horizon, it can be seen that the Poincare section is distorted and finally transits from Kolmogorov-Arnold-Moser (KAM) tori to scattered points. This means that the particle motion becomes more and more chaotic when moving closer to the horizon which is the same as the results of general relativity [13].
@@ 4 Lyapunov exponent of conformal anomaly corrected black hole:
$\frac{e}{m}=-\frac{\sqrt{f_{1}}}{2\psi_{e1}\sqrt{r_{0}-r_{+}}}+\frac{-3f_{2}\psi_{e1}+4f_{1}\psi_{e2}}{4\sqrt{f_{1}}\psi_{e1}^{2}}\sqrt{r_{0}-r_{+}}$ (31)
For the fixed charge to mass ratios, the equilibrium position can be calculated from this relation where
$r_{0}=\frac{2\frac{e}{m}\sqrt{f_{1}}\psi_{e1}^{2}-\sqrt{2}\sqrt{-3f_{1}f_{2}\psi_{e1}^{2}+2\frac{e^{2}}{m^{2}}f_{1}\psi_{e1}^{4}+4f_{1}^{2}\psi_{e1}\psi_{e2}}}{-3f_{2}\psi_{e1}+4f_{1}\psi_{e2}}$ (32)
Moreover,substitute Eq.(4) into the Eq. (26), the Lyapunov exponent $\lambda^{2}$ is modified to be
@@
We next investigate the conformal anomaly correction on the Lyapunov exponent. Now without loss of generality, we choose to switch off the dark radiation $Q=0$ and all the contribution of U(1) charge is given by the electric charge $Q_{e}$ for simplicity. Since the equilibrium position $r_{0}$ will be different for different anomaly correction parameters, we will fix the charge to mass ratio to be $e/m=30,50,100$ respectively and solve $r_{0}$ first by condition Eq.(32). We choose the same black hole parameter as Sec.3 where $M=1$, $Q_{e}=0.5$. Then, by plugging $r_{0}$ into Eq. (35), the relation between Lyapunov exponent $\lambda$ and anomaly central charge $\alpha_{c}$ is explicitly shown in Fig.3.
@@ 5 Conclusion and Discussion:
In this work, we investigate the influence of the conformal anomaly correction on the chaotic particle motions around black hole. For the observational consideration, we only investigate the case for asymptotically flat case. To guarantee that the conclusion holds in more general settings, we choose two different kinds of particle motions to show the influence of conformal anomaly on the chaotic motions. We use two different probes to illustrate the effect of conformal anomaly which is Poincare section and Lyapunov exponent respectively. For Poincare section in fixed $E$ case, we found that by increasing the anomaly parameter $\alpha_{c}$, the original KAM tori in Poincare section is distorted and finally broken into irregular scattered points which means the original integrable motion will transit to the chaotic motions. For Lyapunov exponent, we found that the Lyapunov exponent monotonically increase when we increase the anomaly coefficient which means that conformal anomaly enhances chaos. Therefore, for two different kinds of chaotic motions investigated in Sec.3 and Sec.4, the final results remain the same which is conformal anomaly will enhance chaos near the black hole horizon. This indicates that our conclusion holds in general settings.

### v2

@@ Abstract
The black hole horizon can induce chaotic motion of particles around the black hole. The original integrable motion of particles can transit to the chaotic motion when approaching black hole horizon. In this work, we consider the black hole background where quantum conformal anomaly correction is taken into account. We use Poincare section and Lyapunov exponent as representative probes to illustrate chaos. We investigate the effect of anomaly coefficient on the chaos of particles and for both probes, we find that quantum anomaly generally enhances the chaos of particle motions. As the chaotic orbits of particles can leave an imprint on the gravitational wave signals of the extreme mass ratio inspirals, our results pave the way to detect conformal anomaly effect in actual observations.
@@ 1 Introduction:
Besides the dynamical properties of particles around black hole, black hole itself also has interesting dynamical properties. There are four law of black hole dynamics which is in close analogy with thermodynamical four laws [6]. The discovery of Hawking radiation [7] promotes the four laws of black hole dynamics to black hole thermodynamics and black hole behaves like an ordinary thermodynamical system with its entropy proportional to the horizon area [8]. Moreover, if the cosmological constant can be treated as the thermodynamical pressure [^2] For recent related discussions about this point, see Ref.[9].[10], the equation of state of RN-AdS black hole is the same as the equation of state in van der-Waals liquid-gas system, thus by analogy, the black hole can exhibit rich phase transition structures [11, 12, 13]. For RN-AdS black hole, there exists a critical point where the phase transition becomes second order and the critical exponents coincide with the mean field theory result [11].
@@
The chaotic motion of particles moving around black hole receives lots of attentions  [^4] In this work, we only focus on the particle motions. However, it needs to be stressed that the chaotic motions of extended objects such as strings have also been investigated in Ref.[20, 21] which is useful to understand holographic QCD system., there are many investigations from different point of views. Ref.[22, 23, 24, 25, 26, 27, 28, 29, 30] investigate the chaos bound and its possible violations. Ref.[31, 32, 33, 34] investigates the relation between chaos and black hole phase transition. Ref.[35, 36] investigate if the chaotic motions around black hole can be used to probe the dark matter in the universe. While Ref.[37, 16] considered the chaotic particle motions in modified gravity theories inspired by quantum effects. Finally, recent Ref.[38] considered the imprint of particles’ chaotic motion on the gravitational waves for extreme mass ratio inspirals (EMRI) system.
@@ 3 Poincare section of conformal anomaly corrected black hole:
with the blackening factor corrected by the conformal anomaly given in Eq.(5). By redefining $E\to\frac{E}{m}$ and $p\to\frac{p}{m}$, the mass $m$ can be set to unity without loss of generality.
In order to numerically solve the above equations of motion, we should choose the parameter values without loss of generality. By redefining dimensionless variables $Q\to\frac{Q}{M}$, $r\to\frac{r}{M}$ and $\alpha_{c}\to\frac{\alpha_{c}}{M^{2}}$ in Eq.(5)  [^6] Note that this redefinition is equivalent to choose $M=1$, the dimensionless parameter values are chosen as
@@
The initial values of $r$, $\theta$ and $p_{r}$ are chosen randomly while the initial value of $p_{\theta}$ is solved by the energy constraint Eq.(13). In Fig.1, we show the Poincare section of the particle trajectories which is projected in the $(r,p_{r})$ plane. The section is defined by the constraint condition $\theta=0$ and $p_{\theta}>0$. We gradually increase the energy to see the change of Poincare sections. It can be found that under the influence of conformal anomaly $\alpha_{c}$, by increasing the energy which corresponds to approaching the horizon, it can be seen that the Poincare section is distorted and finally transits from Kolmogorov-Arnold-Moser (KAM) tori to scattered points. This means that the particle motion becomes more and more chaotic when moving closer to the horizon which is the same as the results of general relativity [15].
@@ 4 Lyapunov exponent of conformal anomaly corrected black hole:
$\frac{e}{m}=-\frac{\sqrt{f_{1}}}{2\psi_{e1}\sqrt{r_{0}-r_{h}}}+\frac{-3f_{2}\psi_{e1}+4f_{1}\psi_{e2}}{4\sqrt{f_{1}}\psi_{e1}^{2}}\sqrt{r_{0}-r_{h}}$ (31)
For the fixed charge to mass ratios, the equilibrium position can be calculated from this relation which reads
$x=\frac{2\frac{e}{m}\sqrt{f_{1}}\psi_{e1}^{2}-\sqrt{2}\sqrt{-3f_{1}f_{2}\psi_{e1}^{2}+2\frac{e^{2}}{m^{2}}f_{1}\psi_{e1}^{4}+4f_{1}^{2}\psi_{e1}\psi_{e2}}}{-3f_{2}\psi_{e1}+4f_{1}\psi_{e2}}$ (32)
where $x=\sqrt{r_{0}-r_{h}}$. Moreover,substitute Eq.(4) into the Eq. (26), the Lyapunov exponent $\lambda^{2}$ is modified to be
@@
We next investigate the conformal anomaly correction on the Lyapunov exponent. Now without loss of generality, we choose to switch off the dark radiation $Q=0$ and all the contribution of U(1) charge is given by the electric charge $Q_{e}$ for simplicity. Since the equilibrium position $r_{0}$ will be different for different anomaly correction parameters, we will fix the charge to mass ratio to be $e/m=30,50,100$ respectively and solve $r_{0}$ first by condition Eq.(32). We choose the same black hole parameters as in Eq.(18) where $M=1$, $Q_{e}=0.5$ after non-dimensionalization. Then, by plugging $r_{0}$ into Eq. (35), the relation between Lyapunov exponent $\lambda$ and anomaly central charge $\alpha_{c}$ is explicitly shown in Fig.3.
@@ 5 Conclusion and Discussion:
In this work, we investigate the influence of the conformal anomaly correction on the chaotic particle motions around black hole. For the observational consideration, we only investigate the case for asymptotically flat case. To guarantee that the conclusion holds in more general settings, we choose two different kinds of particle motions to show the influence of conformal anomaly on the chaotic motions. We use two different probes to illustrate the effect of conformal anomaly which is Poincare section and Lyapunov exponent respectively. For Poincare section in fixed $E$ case, we found that by increasing the anomaly parameter $\alpha_{c}$, the original KAM tori in Poincare section is distorted and finally broken into irregular scattered points which means the original integrable motion will transit to the chaotic motions. For Lyapunov exponent analysis, we found that the Lyapunov exponent monotonically increase as we increase the anomaly coefficient which indicates that conformal anomaly enhances chaos. Therefore, for two different kinds of chaotic motions investigated in Sec.3 and Sec.4, the final results remain the same: conformal anomaly will enhance chaos near the black hole horizon. This illustrates that our conclusion holds in general settings.